OpenStudy (anonymous):

Find the following values to the equation of the ellipse (x+5)^2/25 +(y-7)^2/144 = 1. a. Center b. Vertices c. Length of minor axis d. Length of the major axis e. Foci

OpenStudy (anonymous):

here is all you need to answer this question: \[\large \frac {(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 \] where (h,k) is the center vertices for this ellipse: (h,k\(\pm \)b) length of the minor axis: 2a length of major axis: 2b foci: (h,k\(\pm \)c where \(\large c^2 =b^2-a^2\)

OpenStudy (anonymous):

let's find (h,k) first... h= k= ??

OpenStudy (anonymous):

look at your equation and match it up with the one i wrote... what matches with h and k?

OpenStudy (anonymous):

please hurry i have to have this done in 5 minutes or i fail1!

OpenStudy (anonymous):

and thank you! :D

OpenStudy (anonymous):

idk what?

OpenStudy (anonymous):

\(\large \frac {(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\) \(\large \frac {(x+5)^2}{5^2}+\frac{(y-7)^2}{12^2}=1\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so what are the lengths without vvariables? o.O

OpenStudy (anonymous):

for the minor axis, its 2a for the major axis, its 2b you need to tell me what a and b are....

OpenStudy (anonymous):

3 minutes!!!!!

OpenStudy (anonymous):

but what about the ellipse that is nothing but variables?

OpenStudy (anonymous):

a is 5 b is 12

OpenStudy (anonymous):

and am i allowed to have variable in the legths?

OpenStudy (anonymous):

the equation tells you everything you need to know..... you can identify all these quantities i'm asking for just by looking at your equation...

OpenStudy (anonymous):

i already gave you a and b...

OpenStudy (anonymous):

so 2a is 10?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and 2b is 24?>

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the foci and eellipse?

OpenStudy (anonymous):

what is c?

OpenStudy (anonymous):

use the relation i showed u.... in the first post...

OpenStudy (anonymous):

you already know what a and b are.. you can find c...

OpenStudy (anonymous):

hmmm i really do not know this plus i am slightly brain dead btw r u a guy or girl

OpenStudy (anonymous):

dont i just have get the square roote of c^2 and since c^2=119 than the answer would be 10.9? right?

OpenStudy (anonymous):

cookies....why does it matter?

OpenStudy (anonymous):

just wanna know

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

dpanlc?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

idk so i dnt get it wrong

OpenStudy (anonymous):

reaally/?

OpenStudy (anonymous):

well i'm a girl....soooooooo

OpenStudy (anonymous):

yess i dnt wanna be rude kewl glad to have a gal pal